OpenStudy (anonymous):

72:84=_:7 what does the blank equal. how whould u solve it?

ganeshie8 (ganeshie8):

\( a : b = c : d \) is same as \(\large \frac{a}{b} = \frac{c}{d}\) cross multiply \(\large ad = bc\)

OpenStudy (anonymous):

I don't understand what u mean because its 72:84=blank :7 because im doing ratios and I have to simplify it

ganeshie8 (ganeshie8):

\(72:84= x :7 \) is same as \(\large \frac{72}{84} = \frac{x}{7}\) cross multiply \(\large 72 \times 7 = x \times 84\) solve \(x\)

OpenStudy (anonymous):

ohh ok thank you for ur help but I think u would divide it not multiply

ganeshie8 (ganeshie8):

when u cross multiply, division problems turns into a multiplicaiton problem

OpenStudy (anonymous):

ohh

ganeshie8 (ganeshie8):

^ above, solve \(x\) and see wat u get :)

OpenStudy (anonymous):

I got 504

OpenStudy (anonymous):

and I divided 504%7 and I got 72 and I did 48%7 and I got 6 remander 6 so I did 72:48=66:7 right

ganeshie8 (ganeshie8):

\(\large 72 \times 7 = x \times 84\) \(\large \frac{72 \times 7}{84} = x\) \(\large \frac{504}{84} = x\) \(\large 6 = x\)

ganeshie8 (ganeshie8):

so, 72:84=6:7